### v2

# Smaller and More Flexible Cuckoo Filters

Abstract
Cuckoo filters are space-efficient approximate set membership data structures with a controllable false positive rate (FPR) and no false negatives, similar to Bloom filters. In contrast to Bloom filters, Cuckoo filters store multi-bit fingerprints of keys in a hash table using variants of Cuckoo hashing, allowing each fingerprint to be stored at a small number of possible locations. Existing Cuckoo filters use fingerprints of $(k+3)$ bits per key and an additional space overhead factor of at least $1.05$ to achieve an FPR of $2^{-k}$. For $k=10$, this amounts to 1.365kn1.365\,kn bits to store $n$ keys, which is better than 1.443kn1.443\,kn bits for Bloom filters. The +3+3 for the fingerprint size is required to balance out the multiplied FPR caused by searching for the fingerprint at $2^{3}=8$ locations. In the original Cuckoo filter, the number of hash table buckets is restricted to a power of 2, which may lead to much larger space overheads, up to 2.1(1+3/k)kn2.1\,(1+3/k)\,kn bits.
We present two improvements of Cuckoo filters. First, we remove the restriction that the number of buckets must be a power of 2 by using a different placement strategy. Second, we reduce the space overhead factor of Cuckoo filters to 1.06(1+2/k)1.06\,(1+2/k) by using overlapping windows instead of disjoint buckets to maintain the load threshold of the hash table, while reducing the number of alternative slots where any fingerprint may be found. The is parallelized implementation using a producer-consumer method which uses lock-free communication via message buffers.
A detailed evaluation demonstrates that the alternative memory layout based on overlapping windows decreases the size of Cuckoo filters not only in theory, but also in practice. A comparison with other state-of-the art filter types, Prefix filters and Vector Quotient filters (VQFs), shows that the reduced space overhead makes windowed Cuckoo filters the smallest filters supporting online insertions, with similarly fast queries, but longer insertion times. The improved Cuckoo filters are also more flexible in the sense that current implementations of Prefix filters and VQFs only support a restricted set of FPRs.
Our implementation of bucketed and windowed Cuckoo filters, and the workflows to reproduce all results are available at https://gitlab.com/rahmannlab/cuckoo-filters.
keywords:
Cuckoo hashing, approximate set membership, probabilistic filter, power of choices
1 Introduction
Probabilistic filters
are space efficient data structures for approximate fast set membership queries. They have many practical applications, such as for database systems [25], networks [15, 26], storage systems [22] and sequence analysis in computational biology [18, 21, 24]. A probabilistic filter represents a set $K$ of keys from a key universe $\mathcal{U}$ and supports at least two operations: (1) inserting a key, and (2) querying if a key is contained in the set. Instead of storing an exact representation of the set $K$, the keys are fuzzily represented in a way that queries for keys that were inserted into the filter are always correctly answered (no false negative results). Queries about keys that were not inserted are correctly answered most of the time, but sometimes incorrectly positively. The probability of obtaining a false positive answer is called the false positive rate (FPR). The FPR is controlled by the user and related to the space requirements of the filter. To represent an arbitrary set of keys $K$ with cardinality $|K|=n$ with an FPR of $\varepsilon=2^{-k}$, the theoretically optimal space requirement is $nk$ bits. Practical implementations of filters need $Cnk$ bits with an overhead factor $C>1$ that differs between filter types. Different filter types try to achieve small $C$ and fast insertion and query times; often with different trade-offs between space and time efficiency. In addition, some filter variations support additional operations, such as deletion of keys (dynamic filters), counting the number of times a key was inserted (counting filters), or even storing arbitrary values with the keys (probabilistic key-value stores). Other variants optimize the space requirements for a fixed set of keys, where the complete set of keys has to be known before filter construction starts (static filters).
Table 1: Overview of non-extendable filter types supporting online insertions. The overhead factor $C=C(k)>1$ specifies the required space per key; one needs CnkCnk bits to store $n$ keys with an FPR of $2^{-k}$. If a filter only supports a number of slots that is a power of $2$, $C(k)$ can vary between the given formula (best case) and twice that value (worst case; here given as worst-case $C(8)$ for a typical $k=8$ scenario). The number of cache misses for insertion/lookup is a proxy for insertion/lookup time.
Overhead factor $C(k)$, Worst Cache misses for
Name [Ref.] best case (max load) $C(8)$ insert lookup
Bloom [2] 1.443 1.443 $k$ $k$
Standard Cuckoo [12] 1.050(1+3/k)1.050\,(1+3/k) (†) 2.887 1 to $M$ (∗) 2
Improved Cuckoo [this] 1.057(1+2/k)1.057\,(1+2/k) 1.321 1 to $M$ (∗) 2
Quotient [1] 1.053(1+2.125/k)1.053\,(1+2.125/k) (†) 2.665 some (∗∗) some (∗∗)
Vector Quotient [29] 1.075(1+2.914/k)1.075\,(1+2.914/k) (†) 2.933 2 2
Prefix [11] $(1+\gamma)\;(1+2/k)+\gamma/k$ (‡) 1.359 $\geq 1$ (∗∗∗) $\geq 1$ (∗∗∗)
(†): The number of buckets/slots in the filter must be a power of 2.
(‡): The Prefix filter achieves the same FPR for different parameters $\gamma$, here $\gamma=1/\sqrt{2\pi\cdot 25}\;$ [11].
(∗): For Cuckoo filters, $M$ is the maximum random walk length during insertion. Higher values of $M$ achieve lower overhead but require more time.
(∗∗): For the Quotient filter, cache misses depend on the load factor; higher loads (lower overhead) means more cache misses and more time.
(∗∗∗): For the Prefix filter, cache misses depend on the load factor and whether the spare (second level; used if bins in first level are full) is accessed. If the spare is not accessed, insert and queries require 1 cache miss. The number of cache misses in the spare depend on the chosen spare filter (e.g., Cuckoo, Vector Quotient or Blocked Bloom filter).
Filter variants.
Bloom filters [2] were first introduced in 1970 and are still among the most commonly used filters. Bloom filters, when optimally configured for an FPR of $2^{-k}$, store a bit array with $m:=(1/\ln 2)\,nk\approx 1.443\,nk$ bits. Disadvantages of Bloom filters are the relatively high overhead of 44.3% and slow insertion and query times. Blocked Bloom filters [30] achieve faster insertion and query times at the cost of even larger space, and the recent BlowChoc filters [31] keep the advantages of Blocked Bloom filters while reducing the space requirements, sometimes even below those of standard Bloom filters.
However, alternative filter designs often perform better than Bloom filters. Cuckoo filters [12], Morton filters [3], (Vector) Quotient filters [14, 29] and Prefix filters [11] store a $k$-bit fingerprint in a hash table instead of distributed bits in a bit array. This design allows for storing additional values together with the key fingerprints, simply by using additional bits, giving us not only a probabilistic set membership data structure, but also a probabilistic key-value store without additional effort. The FPR is controlled by the fingerprint size, where larger fingerprints lead to smaller FPRs. Different filter types in this class differ in their hash collision resolution strategies. For example, Cuckoo filters resolve collisions using Cuckoo hashing [28, 12] and Vector Quotient filters use ideas from Robin Hood hashing [29]. Depending on the collision resolution strategy, the filters achieve different load thresholds. For a hash table with $s$ slots, the load (or fill rate) $r$ is given by $r:=n/s$, where $n$ is the number of inserted keys. The load threshold (or maximum load factor) is the largest load $r$ for which the filter can be successfully constructed with high probability. To achieve low space overhead we need to design a data structure that has a high load threshold while at the same time limiting the number of possible positions a key’s fingerprint may be stored, as each alternative position would increase the FPR if not compensated for by additional bits per key.
If the whole key set is known in advance, static filters, such as XOR filters [16], (Bumped) Ribbon filters [9, 8] or Binary Fuse filters [17], achieve smaller overhead factors, but are not suitable for streaming applications where the key set $K$ is usually not known in advance.
The above filters, even the non-static ones, require knowledge of the cardinality $n=|K|$ (or an accurate estimate of it) prior to filter construction. In contrast, extendable filters, such as Infini filters [5], Aleph filters [6], dynamic Cuckoo filters [4] or consistent Cuckoo filters [23] increase their capacity if more keys arrive than originally planned for, at the cost of a higher memory usage and/or higher FPR. Adaptive Cuckoo filters [27] can react to false positive queries by re-arranging fingerprints in the Cuckoo filter. To recognize a false positive, adaptive Cuckoo filters require the whole key set to be additionally stored in a hash table, and are thus only practical in applications where the Cuckoo filter only acts as a fast pre-filter.
Figure 1: (A) Best-case overhead factors for different filters and FPRs. The standard Cuckoo filter is the bucketed (2,4) Cuckoo filter (orange). (B) Number of required buckets in the original Cuckoo filter [12] (dashed blue staircase) vs. an optimally sized Cuckoo filter (orange diagonal).
In this work, we focus on fingerprint based non-extendable filters supporting online insertions; an overview is given in Table 1. In contrast to the Bloom filter, the (best possible) overhead factor of the fingerprint based filters decreases with increasing $k$ (see also Figure 1), but many of these filters have additional technical or efficiency-related restrictions, which can increase the overhead by a factor of (almost) 2, and hinders flexible use.
Contributions and Outline.
We improve upon Cuckoo filters in two ways: (1) We remove technical restrictions that until now only yield a small space overhead under certain conditions on the number $n$ of inserted keys. (2) We change the memory layout to overlapping windows instead of non-overlapping buckets, obtaining higher hash table load with fewer possible locations for each key, thus needing fewer bits of memory overall. The advantages of the windowed memory layout also apply to other Cuckoo filter variants, such as the dynamic Cuckoo filter [4] or the adaptive Cuckoo filter [27], but we limit our evaluation to the standard Cuckoo filter with fixed capacity. Our implementation is parallelized using independent subfilters which can be filled by separate threads without locking or complex communication. We obtain universally usable filters that are both small and fast.
After providing background on Cuckoo hashing and Cuckoo filters in Section 2, we describe how to improve Cuckoo filters in Section 3 and provide implementation details in Section 4. In Section 5, we provide a detailed evaluation of bucketed and windowed Cuckoo filters and compare them to other state-of-the-art filters. Section 6 concludes.
2 Background
2.1 Multi-way Cuckoo hashing
In standard Cuckoo hashing [28], a set of keys $K=\{x_{1},x_{2},\dots,x_{n}\}\subset\mathcal{U}$ is stored (exactly) in a hash table with $s\geq n$ slots; indexed $0,1,2,\dots,s-1$. After inserting $n$ keys, the fill rate (or load factor) of the hash table is $r:=n/s$. The hash table may store additional data associated with each key. Each key $x$ may be inserted into one of two possible slots, computed using two hash functions $f_{1}:\mathcal{U}\to[s]:=\{0,\dots,s-1\}$ and $f_{2}:\mathcal{U}\to[s]$, randomly chosen from a universal family. If both slots $f_{1}(x)$, $f_{2}(x)$ for key $x$ are already occupied, one of the keys at $f_{1}(x)$ or $f_{2}(x)$ is removed and re-inserted into its alternative slot. This slot might also be occupied, so a chain of removals and re-insertions starts until a free slot is found. This can lead to long walks or end in a cycle, and the maximum load factor that can be achieved with a high probability in this setting is $1/2$, a rather low value.
Table 2: Rounded theoretical load thresholds for $(2,l)$ Cuckoo hashing [32].
layout $l=1$ 2 3 4
buckets (2-ary) 0.5 0.8970118682 0.9591542686 0.9803697743
windows (2-ary) 0.5 0.9649949234 0.9944227538 0.9989515932
To achieve higher loads and consequently use less memory, Cuckoo hashing has been generalized in the following three ways.
  1. 1.
    In $d$-ary Cuckoo hashing, $d$ independent hash functions $f_{1},f_{2},\dots,f_{d}:\mathcal{U}\to[s]$ are used to compute $d$ candidate positions for each key [13].
  2. 2.
    In $(d,l)$ bucketed Cuckoo hashing, the hash table is divided into $B=\lceil s/l\rceil$ non-overlapping buckets, each containing $l$ slots. The $d$ hash functions $f_{1},\dots,f_{d}:\mathcal{U}\to[B]$ pick alternative buckets. Each key can be inserted into any of the $l$ slots inside the $d$ buckets [7, 33].
  3. 3.
    $(d,l)$ windowed Cuckoo hashing works similarly to $(d,l)$ bucketed Cuckoo hashing, but uses overlapping windows instead of non-overlapping buckets [20]. Each key can be inserted into one of $dl$ slots, but each window overlaps by l1l-1 positions with its surrounding windows, hence the total number of windows is W=sl+1W=s-l+1.
Theoretical asymptotic load thresholds for these generalizations have recently been computed by Stefan Walzer [32]; see Table 2. The load threshold increases with $d$ and $l$ and is higher for windows compared to buckets. However, larger $d$ leads to increased query times, since we search a key in each of the $d$ buckets or windows at different memory locations, each likely causing a cache miss, while a search inside a bucket or window is local. Hence, we consider only $d=2$. Since existing Cuckoo filter implementations use buckets, it is natural to ask whether a windowed memory layout reduces the memory overhead while retaining the fast query times of Cuckoo filters also in practice.
2.2 Cuckoo filters
A Cuckoo filter is a probabilistic data structure that is based on Cuckoo hashing, introduced using $(2,4)$ bucketed Cuckoo hashing [12]. A Cuckoo filter uses a fingerprint hash function $f_{0}:\mathcal{U}\to[2^{q}]$ that computes a $q$-bit fingerprint for a key. Each key is assigned to two distinct buckets $b_{1}$ and $b_{2}$ from a total of $B$ buckets. A bucket consists of a constant number of available slots (4 in [12]), and each slot may hold a fingerprint. Therefore, the fingerprint of a key can be stored in any of $2\cdot 4=8$ slots. In order to guarantee an FPR of $2^{-k}$, with 8 possible locations for each key, the fingerprint size needs to increase to $q:=k+3$ bits, yielding an FPR bounded by $8\cdot 2^{-(k+3)}=2^{-k}$.
Since only store the fingerprint is stored, the two bucket addresses must be restricted so that one can be calculated from the other and the fingerprint. If we do not want to sacrifice a bit that remembers whether the fingerprint is stored in its first or second bucket, we must use a symmetric function to compute both buckets. The standard Cuckoo filter solves this problem as follows [12]. A hash function $f_{1}:\mathcal{U}\to[B]$ computes the first bucket address of a key $x$, given by $b_{1}=f_{1}(x)$. The second bucket address is the bit-wise XOR of a hash of the fingerprint $h(f_{0}(x))$ and the first bucket address $b_{1}$, where $h:[2^{q}]\to[B]$. Conversely, the first bucket address is obtained back by XORing $b_{2}$ with the same fingerprint hash. The XOR operation is only guaranteed to return a valid address if the number of buckets is a power of 2. A proposed simplification [10] uses the $q$-bit fingerprint $f_{0}(x)$ directly instead of a hash value $h(f_{0}(x))$.
Insertions.
Inserting a key $x$ with 2 hash functions and a bucket size of $l$ works as follows. If any of the $2l$ slots where the fingerprint of $x$ can be stored is empty, it is stored there, giving priority to slots in the first bucket $b_{1}$. If all $2l$ slots are full, pick a random one of these slots, remove the stored fingerprint, say of key $x^{\prime}$, and insert fingerprint $f_{0}(x)$ at the now free slot. Then, attempt to insert the removed fingerprint $f_{0}(x^{\prime})$ into its alternative bucket (which can be computed from its former address and the fingerprint itself, without knowledge of $x^{\prime}$). This procedure is repeated for up to a given number of steps. The insertion fails if no free slot can be found. In this case, the filter size must be increased and all insertions repeated. The probability of failure approaches zero if the number of buckets is large enough (and enough steps are allowed). In the described $(2,4)$ configuration, $B=1.02\cdot n/4$ buckets, or 1.02n1.02\,n slots are sufficient in theory to insert all $n$ elements. In practice, in order to limit the length of the random insertion walks, the number of buckets and slots is chosen larger, and the original work [12] uses 1.05n1.05\,n slots. It has to be noted that a Cuckoo filter (if not over-provisioned) cannot tolerate many additional insertions; these will simply fail.
Queries.
To query the presence of a key in a $(2,4)$ bucketed Cuckoo filter, search all eight possible slots and return True if and only if the fingerprint $f_{0}(x)$ was found at any of the slots. Queries are fast, because at most two distinct memory locations (buckets $b_{1}$ and $b_{2}$) need to be accessed; the remaining memory accesses are local within each bucket.
3 Flexible windowed Cuckoo filters
We improve upon the standard Cuckoo filter in two ways: First, we introduce a different way to find the alternative bucket (or window) from the given one and the fingerprint, using a signed offset. Second, we use overlapping windows instead of disjoint buckets.
3.1 Moving between alternative buckets or windows
In the standard Cuckoo filter [12], there are two alternative buckets to store the $q$-bit fingerprint $f_{0}(x)$ of a key $x$. If $b$ is one of the buckets, $b^{\prime}$ is the alternative bucket, and $h:[2^{q}]\to[B]$ is a hash function, then setting $b^{\prime}=b\oplus h(f_{0}(x))$ is symmetric in $b,b^{\prime}$ but requires that the number of buckets satisfies $B=2^{\beta}$ for some integer $\beta$ to guarantee valid bucket addresses $b,b^{\prime}\in[B]$. This may introduce a significant memory overhead (up to an additional factor of 2.0; see Figure 1B). An alternative [16] proposed $b^{\prime}:=(B-(b+f_{0}(x)))\bmod B$, which is also symmetric in $b,b^{\prime}$, but has no restrictions on $B$. However, it does not allow further randomization, which makes it vulnerable to adversarial input data and introduces strong anti-correlation between the locations of $b$ and $b^{\prime}$.
We use a different approach, consisting of the first bucket hash function $f_{1}:\mathcal{U}\to[B]$ and an offset hash function $f_{2}:[2^{q}]\to[B-1]$, mapping any fingerprint to an offset less than $B$. For a key $x$, we have the first bucket address $b=f_{1}(x)$ and $b^{\prime}:=(b+f_{2}(f_{0}(x)))\bmod B$ with $b^{\prime}\neq b$. From $b^{\prime}$, bucket $b$ cannot be reconstructed symmetrically, but asymmetrically as $b=(b^{\prime}-f_{2}(f_{0}(x)))\bmod B$. Hence, we need to use a bit in addition to the fingerprint to store the current bucket choice ($b$ or $b^{\prime}$) that tells us whether we need to add or subtract the offset to obtain the alternative bucket.
It seems disadvantageous that we need one extra bit in each slot. However, when we query the presence of a key, we only consider the key to be present if both the fingerprint and the choice bit are identical. Therefore, we can use the choice bit as one of the 3 extra bits needed to counteract the 8 possible slots. We do not require that this bit is uniformly distributed between 0 and 1 across the filter: Assume that fraction $p$ of all fingerprints (of length $q=k+2$) are stored in their respective first bucket with choice bit 0, and the remaining fraction 1p1-p in their second bucket with choice bit 1. A false positive occurs if we find the fingerprint together with choice bit 0 in the first bucket (probability $\leq 4\cdot 2^{-q}\cdot p$) or with choice bit 1 in the second bucket (probability $\leq 4\cdot 2^{-q}\cdot(1-p)$). So the total probability is bounded by $4\cdot 2^{-q}\cdot(p+1-p)=2^{-k}$, independently of $p$.
3.2 Windowed layout
Figure 2: Inserting a key’s fingerprint into a Cuckoo filter with window size 4 and an FPR of $2^{-5}$: Compute both window addresses and search for an empty slot in either window. If both windows are full, as shown here, remove a random fingerprint from one of the windows (here, the orange fingerprint), compute its two possible windows based on the choice and offset bits and try to re-insert the fingerprint at one of the alternative slots. Here, the removed fingerprint can directly be re-inserted into its current window, since the last position in the window is still empty.
The standard Cuckoo filter uses (2,4) bucketed Cuckoo hashing [12]. Alternatively, the array may be divided into $W$ overlapping windows of size $l$, where each window overlaps with the previous window by l1l-1 slots. Using windows has been shown to increase the load threshold for Cuckoo hash tables in comparison to buckets [32, 20]; see Table 2. This allows us to move from $(2,4)$ bucketed Cuckoo hashing to $(2,2)$ windowed Cuckoo hashing while maintaining a comparable load threshold.
However, new complications arise: Any slot of the hash table now belongs to $l$ different windows, and from the location alone we do not know the window number. Hence, we also need to store the offset from the window start, a number in $[l]$, taking $\log_{2}l$ bits. (For convenience, $l$ should therefore be a power of 2, such as 2 or 4.) Fortunately, again, these need not be extra bits, but can be used as part of the bits that counteract the FPR multiplier. For concreteness, in a $(2,2)$ windowed Cuckoo filter, we use a $k$-bit fingerprint, one additional choice bit to indicate which of the two alternative window locations we are using, and one additional offset bit within the window (first or second slot). Similarly to the choice bit, the distribution of the window offset bits need not be uniform over all $2^{l}$ possible values.
Inserting or querying a key works analogously to the version with buckets. To insert a key, check if any of the $2l$ slots is empty. If yes, insert the fingerprint into the empty slot. Otherwise, remove a random fingerprint and try to re-insert it into an alternative slot. To re-insert the removed fingerprint, compute both its current and alternative window based on the stored bits (Figure 2). Run this loop of removal and re-insertion until an empty slot is found, but at most for a fixed number of steps. For queries, return True if and only if any of the $2l$ possible slots contain the correct fingerprint and the correct choice and window offset bits for the currently searched slot.
Thus, the main differences of our proposal to the previous Cuckoo filter [12] are that here,
  1. 1.
    the additional bits are not (randomized) extensions of the fingerprint, but have meaning as choice bit and offset bits,
  2. 2.
    we reduce the number of offset bits from 2 to 1, saving space, because the load threshold for Cuckoo hashing with two slots in windows and for four slots in buckets are comparable ($0.965$ vs. $0.980$). More precisely, filling the hash tables slightly below their theoretical load thresholds, the overhead factor is reduced from $1.05(1+3/k)$ in [12] to $1.06(1+2/k)$, e.g., from $1.365$ to $1.272$ for $k=10$ (an FPR of $1/1024$).
4 Implementation Details
We now discuss our implementation using optimized just-in-time compiled Python, how we distinguish empty from full slots in the hash table, how we execute bit-parallel queries over several slots, and how we parallelize insertion. Further aspects, such as the concretely used hash functions, are discussed in Appendix A.
4.1 Just-in-time compilation
We have implemented the improved Cuckoo filters for integer keys in just-in-time compiled Python using the numba package [19] with typed numpy arrays. This offers several benefits, such as highly optimized machine code, the use of LLVM intrinsics, and the option to choose parameters during runtime, but before compilation. For example, this allows us to provide all parameters of the hash functions as compile-time constants for additional optimizations, achieving speeds comparable to compiled C or C++ code, or sometimes even faster, due to the increased possibilities for optimization. Since most state-of-the-art filters are implemented in compiled languages, we compared our reference implementation of the original Cuckoo filter with an existing C++ implementation [12]; see Appendix D. Since both implementations achieve similar running times, using just-in-time compiled Python comes with little penalties and allows us to compare our implementation with other state-of-the-art filters written in low-level languages. The code is available at https://gitlab.com/rahmannlab/cuckoo-filters.
4.2 Distinguishing full slots from empty slots
We have so far not discussed how we can distinguish an empty slot from a full slot. The hash function $f_{0}:\mathcal{U}\to[2^{k}]$ maps a key $x$ to its $k$-bit fingerprint $f_{0}(x)$. We use one of the possible fingerprint values (by convention 0) to indicate an empty slot. Hence, we reduce the set of valid fingerprints and instead use a hash function $f^{\prime}_{0}:\mathcal{U}\to[2^{k}-1]$. The fingerprint is given by $f_{0}(x):=f^{\prime}_{0}(x)+1$. The loss of one fingerprint value leads to a higher FPR of $1/(2^{k}-1)$, the probability that two random fingerprints collide. Although this effect is measurable for small $k$, it can be considered as negligible for typically used larger values of $k$; see also Figure 4 in Section 5.2. The same compromise was made for the original Cuckoo filter [12].
4.3 Bit-parallel engineering
We adapt the bit-parallel optimizations from Fan et al. [12] for windowed Cuckoo filters for small enough $k$, where several slots fit into a single 64-bit integer. We discuss them by example using $(2,4)$ windowed Cuckoo filters with $k=5$, which has $q:=k+3=8$-bit slots. For other configurations, similar ideas are implemented with appropriate modifications. For too large $k$, we examine each slot separately at the cost of throughput.
We load the contents of a window into a 64-bit integer register that we (conceptually) partition into four adjacent slots (and some unused bits). To check whether there is an empty slot (all zeros), we use two pre-computed masks, lo and hi = lo << (q-1), indicating the least significant and most significant bits in each slot, respectively. By computing the bit pattern e := (x - lo) & (˜x) & hi, the 1-bits of e are those 1-bits of hi that indicate which slots are empty. If e is non-zero, there exists an empty slot and its index (from the right) can be computed by dividing the number of trailing zeros of e (obtained by a cttz CPU instruction) by q. In the example , the result is nonzero and has 7 trailing zeros, indicating that slot 0 in $x$ is empty. Comparing all 4 slots at the same time in this bit-parallel manner saves the overhead of a loop or separate comparisons.
                slot        3 |        2 |        1 |        0
                hi = 10000000 | 10000000 | 10000000 | 10000000   high bits in each slot
                 x = 10001111 | 00000000 | 11100110 | 00000000   window contents x
                lo = 00000001 | 00000001 | 00000001 | 00000001   low bits in each slot
            x - lo = 10001101 | 11111111 | 11100100 | 11111111   -: subtraction
                ~x = 01110000 | 11111111 | 00011001 | 11111111   ~: bit-wise inversion
       (x-lo) & ~x = 00000000 | 11111111 | 00000000 | 11111111   &: bit-wise and
  (x-lo) & ~x & hi = 00000000 | 10000000 | 00000000 | 10000000   result e, 7 trailing 0s
Similarly, to search for a fingerprint in any of the four slots, we can create a single bit mask that contains the valid bit patterns for each slot. For $(2,4)$ windowed Cuckoo filters, we use one choice bit and two window offset bits. Assume that the 5-bit fingerprint is 11100, that we are examining window choice 1, and that the offset bits increase with decreasing slot number. Then the valid slot bit patterns can be encoded in a mask y, and we can check if the bit patterns agree in one of the slots (and which one) by taking the bit-wise XOR xˆy and checking for a zero slot like above. In the following example, we find the fingerprint in slot 1.
      slot        3 |        2 |        1 |        0
  bit type CWWFFFFF | CWWFFFFF | CWWFFFFF | CWWFFFFF   Choice/Window offset/Fingerprint
       y = 10011100 | 10111100 | 11011100 | 11111100   valid bit pattern per slot
       x = 10001111 | 00000000 | 11011100 | 00000000   window contents x
   x ^ y = 00010011 | 10111100 | 00000000 | 11111100   ^: XOR
4.4 Parallelization
We parallelize the filter as follows: We create $F$ independent subfilters, such that the total number of buckets (or windows) $B$ is divided equally among the $F$ subfilters, with $\lceil B/F\rceil$ buckets or windows per subfilter. Each subfilter is a cache-line-aligned bit array. For inserting or querying a key, the subfilter index is computed by an additional hash function $f_{\text{sub}}:\mathcal{U}\mapsto[F]$. The bucket (or window) address hash function now maps the key $x$ to any of the $\lceil B/F\rceil$ buckets (or windows) in the subfilter. To insert keys into the subfilter, one main thread delegates the work to one thread per subfilter, responsible for performing the subfilter insertion. To avoid locks, the threads communicate via message buffers using a consumer-producer principle, as in [34]. The main thread computes the subfilter number $f_{\text{sub}}(x)$ for each key $x$ and copies the key to a buffer assigned to this subfilter. For each subfilter, several such buffers exist, and these buffers are only read by one consumer thread responsible for insertions into this particular subfilter. If a buffer is full, the main thread sets an atomic volatile ready_to_read flag and continues inserting keys into other buffers. Among the buffers assigned to a specific subfilter, the corresponding consumer thread searches for a buffer marked as ready_to_read and inserts the keys into the subfilter. After all keys in the buffer are inserted, the consumer thread marks the buffer as ready_to_write indicating to the main thread that it can again insert keys into the buffer.
Assuming that queries are processed once all keys are inserted, and given that queries only need read access, it follows that no synchronization between query threads is needed. Queries are thus distributed evenly across all threads.
5 Evaluation
We start by comparing bucketed with windowed Cuckoo filters. We first evaluate the time-memory trade-off for increasing loads (Section 5.1). In Section 5.2, we evaluate the actual FPRs, loads and overhead factors for different desired FPRs at target loads of $0.98\cdot T$, where $T$ is the theoretical load threshold for a given configuration (Table 2). More technical aspects, such as achievable loads, depending on the maximum random walk length during insertion or on the number of inserted keys, are given in Appendix B. Performance benchmarks of our implementation are shown in Section 5.3.
Then, we compare the throughput of our Cuckoo filter implementations with two existing implementations of state-of-the art filters, the Vector Quotient filter and the Prefix filter (Section 5.4). We omit static filters in order to focus on filters used in applications where the whole key set is not known in advance, but where we have a good estimate of its size, which is typical in genome research applications in bioinformatics. In addition, we exclude other filters, like Bloom filters, Morton filters or Quotient filters that showed worse results compared to Prefix and Vector Quotient filters in previous analyses [11, 29].
Vector Quotient filter (VQF) and Prefix filter.
We evaluate the original Vector Quotient filter implementation from [29] (downloaded from https://github.com/splatlab/vqf, written in C++) and the original Prefix filter implementation [11] (from https://github.com/TomerEven/Prefix-Filter, written in C++) using a Cuckoo filter or VQF as a spare.
Experimental setup.
Evaluations were run on a PC workstation with an AMD Ryzen 9 9950X CPU with 16 cores and hyperthreading and 64 GB of DDR5 memory (6000 MHz, CL40). All benchmarks were performed on random unsigned 64-bit integer keys, generated with numpy. Reported times are wall times, including just-in-time compilation and excluding data load time, except for multi-threaded insertion, where reader and inserter threads load and insert keys in parallel. We report averages over five runs.
For insertions, we perform lookup-and-insert operations, i.e., we only insert a key if it is not already classified as present in the filter, except when comparing the throughput between filters in Section 5.4.
5.1 Time-memory trade-off
Figure 3: Time (left) and memory (right) requirements for inserting a billion ($10^{9}$) keys into different Cuckoo filter types (distinguished by color) at different loads (x-axis) for a fixed FPR of $2^{-10}$ and a fixed maximum random walk length of 10 000. Dashed vertical lines correspond to load thresholds. Memory use for windows and buckets coincides, i.e. below their respective thresholds, the yellow and green lines overlap, as well as the blue and orange lines.
Figure 3 shows the time and memory requirements for inserting a billion keys with a fixed FPR of $2^{-10}$ at different final load factors. Higher loads result in higher insertion times, with steep increases as the load threshold is approached. For low loads of 0.8, insertion times are comparable. At load 0.9, already $(2,2)$ buckets are impossible and the time for $(2,2)$ windows has increased from 85 seconds to 140 seconds, but the time for both $(2,4)$ configurations has only slightly increased from 65 seconds to 94 seconds at most. Windowed Cuckoo filters have higher load thresholds compared to bucketed Cuckoo filters with the same number of slots per bucket/window. Therefore, at the same load, the windowed $(d,l)$ Cuckoo filter is faster than the bucketed version.
The memory requirements depend on the window size or bucket size and the load, but there is no difference between windowed and bucketed versions. At the same load, Cuckoo filters with a window size or bucket size of 2 are smaller compared to Cuckoo filters with a size of 4 because one less bit per slot is needed. Windowed Cuckoo filters achieve higher loads and faster insertion times than their bucketed counterparts at fixed load, so they are an overall improvement.
5.2 Actual overhead factors, loads and FPRs
Figure 4: Comparison of actual properties (memory overhead, actual load, empirical FPR) of Cuckoo filter types (distinguished by color) designed for FPRs of $2^{-k}$ for varying $k$ (x-axis), using a fixed maximum random walk length of 10 000, with $n=2\times 10^{9}$ keys and 5 subfilters for parallelization. Insertions are performed only if the filter does not already report a key as present. Left: Actual memory overhead factor $C$ for a Cuckoo filter with CnkCnk bits for a target FPR of $2^{-k}$; lower is better. Middle: Empirical loads. Right: Empirical FPRs, relative to the target FPR of $2^{-k}$, as log ratios. Points below the horizontal red line at 0.00 have an actual FPR that is better than the target FPR.
According to the observations of Appendix B, we choose a filter size that ensures a load of at most 98% of the theoretical load threshold and a maximum walk length of 10 000 steps. However, the actual maximum random walk length is usually much lower; see Appendix C. Using $n=2\cdot 10^{9}$ random 64-bit integer keys, we now compare the resulting actual memory overhead factors, empirical FPRs and observed load factors for different values of $k$ (Figure 4). The actual memory overhead factor is obtained by dividing the actual number of used bits by $nk$. The actual load is the number of occupied slots, divided by the number of total slots. The empirical FPR is obtained by querying $n^{\prime}=2\cdot 10^{9}$ random keys that were not previously inserted into the filter, computed as the number of times the lookup erroneously returned True, divided by $n^{\prime}$.
Overhead factors.
For all Cuckoo filter variants, the overhead factor decreases for increasing $k$ (i.e., for lower FPRs; see Table 1 and Figure 4 left). Windowed Cuckoo filters with window size 2 have the lowest overhead factors, since a window size of 2 saves one bit per slot compared to buckets or windows of size 4, and they achieve high load thresholds. Although buckets of size 2 use the same number of bits per slot, their load threshold is much lower, and they therefore have higher overhead factors compared to windows of size 2. For small $k$, buckets of size 2 have a lower overhead factor compared to windows of size 4, but from $k\geq 8$, the overhead factor for windows of size 4 is lower, since the higher load compensates for the additional bit per slot. Cuckoo filters with buckets of size $4$ have higher overhead factors compared to windows of size $4$ due to the higher load threshold.
Empirical loads.
As shown in Figure 4 (middle), for $k\geq 8$, the empirical load is independent of $k$ and corresponds to the targeted load of 98% of the load threshold. For smaller $k$, the observed load is actually lower. Since small fingerprints have high collision probabilities (of $1/(2^{k}-1)$), several keys are reported as already present during the performed lookup-and-insert operations and not inserted again, causing an overall lower load.
False positive rates.
We target an FPR of $2^{-k}$ by using $k+2$ or $k+3$ bits per slot, for a window or bucket size of 2 or 4, respectively. In practice, several other effects increase or decrease the actual FPR. Figure 4 (right) shows log-ratios between empirical FPR and $2^{-k}$. For $k\geq 5$, the observed FPR is lower than $2^{-k}$, because the filter is not full. Actual non-present keys have a small chance of hitting an empty slot instead of a stored fingerprint, decreasing the FPR. Since the load decreases for small $k$ (see above), the observed FPR should also decrease for small $k$. However, for $k\leq 4$, the window-based filters show a comparably high FPR. Here we see the effect of sacrificing one fingerprint value to indicate empty slots, e.g. for $k=3$, we do not have 8 but only 7 usable fingerprints. Since for the windowed versions, the fingerprint size is only $k$, this effect is severe. For the bucketed versions, we have 1 or 2 additional fingerprint bits (for bucket sizes 2 or 4, respectively) and no window offset bits, so the effect is less severe. While the interplay of these different effects on the FPR for small $k$ is interesting, it is not so important in practice, as most applications aim at FPRs of $2^{-k}$ with $k\gg 5$.
5.3 Performance Benchmarks
Figure 5: Performance benchmarks for (A) insertions, (B) successful and (C) unsuccessful lookups of $2\cdot 10^{9}$ integer keys, with a maximum random walk length of 10 000, averaged over five runs. Top: Throughput in million keys per second for different values of $k$ using five subfilters (inserter threads) or five query threads; higher throughput is better. Bottom: Speedup factor for varying number of threads, for a fixed FPR of $2^{-10}$.
We evaluate insertion throughput and query throughput separately for successful and unsuccessful queries for different $k$, with a load factor of 98% of the load threshold. Figure 5 (top) shows the throughput in million keys per second (wall time) using five inserter and query threads and different values of $k$. Insertion throughput is lower compared to query throughput. This is expected since the random walk may cause several cache misses. In contrast, queries incur at most 2 cache misses. The $(2,4)$ windowed Cuckoo filters have the highest insertion throughput, and $(2,2)$ windowed Cuckoo filters have the lowest insertion throughput, which is in concordance with the results from Section 5.1. The throughput of successful and unsuccessful queries is similarly high (Figure 5B and 5C).
For $l=4$, the whole bucket or window fits into one 64-bit integer for $k\leq 13$. In this case, the optimizations described in Section 4.3 are applied. For $l=4$ and $k\geq 14$, we check each slot in the bucket or window separately, hence query throughput drops. For $l=2$, both windows (or buckets) fit into one 64-bit integer for $k\leq 14$, and throughput drops for $k\geq 15$. The increased throughput at $k\in\{5,13\}$ for $l=4$ and at $k\in\{6,14\}$ for $l=2$, is obtained due to the optimized memory access when the number of bits in a slot is a multiple of 8. Small variations for different values of $k$ may be caused by different compiler optimizations during just-in-time compilation that can be applied to individual fixed values of $k$.
The speedup factor from parallelization is almost linear, both for insertions and for queries for up to 5 threads (Figure 5 bottom, see Appendix 4.4 for details). For more threads, the distributing main thread starts to become the bottleneck. The speedup for insertions is slightly lower than for queries, probably due to higher memory bandwidth utilization.
5.4 Comparison with Prefix Filters and Vector Quotient Filters
Table 3: Throughput in million keys per second (single threaded) and memory comparison of bucketed (b) and windowed (w) Cuckoo, Prefix and Vector Quotient filters. All filters are configured such that they have $2^{30}$ slots and the load is selected such that all insertion succeed with a probability of $\approx 1$ ($T\cdot 0.98$ for Cuckoo, $1.0$ for Prefix and $0.92$ for Vector Quotient filters).
Target FPR $2^{-8}$ Target FPR $2^{-13}$ Target FPR $2^{-14}$
Throughput Throughput Throughput
Filter insert lookup $C$ insert lookup $C$ insert lookup $C$
Cuckoo (b, 2) 4.11 42.88 1.39 4.19 41.85 1.29 4.34 51.67 1.28
Cuckoo (b, 4) 6.06 36.27 1.42 6.43 46.61 1.28 5.81 22.25 1.26
Cuckoo (w, 2) 3.04 30.01 1.31 3.02 28.07 1.21 3.12 35.26 1.20
Cuckoo (w, 4) 7.83 28.45 1.40 8.56 33.16 1.25 7.35 18.41 1.24
Prefix (VQF) 31.32 43.81 1.44
Prefix (CF12) 30.16 44.60 1.45
Vector Quotient 21.59 27.76 1.48
Before comparing the throughput of our Cuckoo filter implementation with other state-of-the-art filters, we note several constraints that complicate a fair comparison.
  1. 1.
    The Vector Quotient filter (VQF) only works if the number of slots is a power of 2. Hence, we set the number of slots to $2^{30}$. In practice, the space overhead could thus be considerably larger for the VQF. In contrast, our Cuckoo filter implementation and the Prefix filter are flexible concerning the number of slots.
  2. 2.
    Both the VQF and the Prefix filter only support a FPR of $\approx 2^{-8}$ (and $2^{-16}$ for the VQF), and their performance is optimized for this special case. A comparable Cuckoo filter with 8 bits per slot has an FPR of $2^{-5}$. In contrast, we support all FPRs with different optimizations due to the advantages of just-in-time compilation, but at the cost of being less optimized for one special case.
  3. 3.
    The available Prefix filter implementation only works on CPUs that support AVX512 instructions; there is no fallback implementation.
Throughput.
We evaluated the insertion and lookup throughput (50% contained and 50% not contained keys) on random 64-bit integers (see Table 3). Prefix filters have the highest insertion and lookup throughput. Cuckoo filters have the lowest insertions throughput due to the many cache misses compared to $\approx 1$ cache miss for Prefix filters and 2 for Vector Quotient filters. For lookups, Vector Quotient filters and Cuckoo filters have similar throughput due to the same amount of cache misses. For $k=8$ Prefix filters have the highest throughput, but in the optimized cases (i.e., Cuckoo filters with 16 bits per slot), bucketed Cuckoo filters have a higher lookup throughput compared to Prefix and VQF filters.
Space overhead.
The space overhead depends on the FPR and a valid comparison is thus only possible for the same FPR. For a given FPR, the $(2,2)$ windowed Cuckoo filter has the smallest space overhead compared to Prefix, VQF and bucketed Cuckoo filters (see Table 3).
Summary.
Our design has a smaller space overhead compared to state-of-the are filters, while lookup times remain competitive. If fast insertion times are essential, and an FPR of $2^{-8}$ is tolerable, Prefix filters are preferable.
6 Conclusion and Discussion
We introduced improvements of Cuckoo filters, enabling their more flexible use and reducing their overhead factor from $1.05(1+3/k)$ to $1.06(1+2/k)$, switching from a $(2,4)$ bucketed layout to a $(2,2)$ windowed layout. This is a practical improvement for relevant FPRs. For an FPR of $2^{-8}$, $(2,2)$ windowed Cuckoo filters need $\approx 10.6$ GB to index 10 billion keys, compared to $\approx 11.6$ GB for $(2,4)$ bucketed Cuckoo filters. Another interesting aspect of this work is that just-in-time compiled Python code can adjust to run-time specified parameters (such as $k$), resulting in a flexible implementation of Cuckoo filters that is competitive in speed with a C++ implementation.
In comparison to other state-of-the-art filters, our implementation of $(2,2)$ windowed Cuckoo filters has the smallest space overhead and competitively fast lookup times, especially for optimized cases, such as $k=14$, where four 16-bit slots can be examined in a bit-parallel manner in a single 64-bit integer. In that case, the overhead factor of $1.20$ even outperforms static XOR filters [16] that need a factor of $1.23$, originally advertised as “faster and smaller than Bloom and Cuckoo filters”. (Of course, by now, better static filters exist [9].)
For $k=8$, Prefix filters are highly optimized concerning fast insertions, here Cuckoo filters are not competitive in terms of speed. In this case, the increased speed may outweigh the slightly larger space overhead of Prefix filters. For even lower values of $k$, Blocked Bloom filters are simpler and can be competitive.
In summary, the low space requirements and fast query times make the improved Cuckoo filters valuable candidates for many applications, e.g., in DNA sequence analysis applications [33], where large exact hash tables may be replaced by smaller filters at the cost of a few false positive queries.
Appendix
Appendix A Hash functions
The hash functions for addressing the subfilter, the buckets or windows, and for computing the fingerprints do not need to satisfy particularly strong properties, but they need to be reasonably fast to compute and have sufficiently many free parameters that can be chosen. In our implementation, we expect $u$-bit unsigned integer keys $x$ with $u\leq 64$ and use linear hash functions of the form $x\mapsto(ax)\bmod b$, where $a$ is a randomized integer, and $a$ and $b$ are coprime. Note that there is always an implicit mod $2^{64}$ operation after every operation due to the unsigned 64-bit arithmetic; so in fact, we are computing $x\mapsto((ax)\bmod 2^{64})\bmod b$. Modifications for a more uniform distribution of the hash values are applied if $b$ is divisible by 2 or a power of 2: If $b$ is a power of 2, we only consider the $\log_{2}b$ most significant bits of $ax$ and use $x\mapsto((ax)\gg(u-\log_{2}b))\bmod b$. If $b$ is even but not a power of 2, we perform a bitwise XOR operation between the $q:=\lfloor u/2\rfloor$ most significant bits and the $u/2$ least significant bits of $(ax)$ first, resulting in $x\mapsto((ax)\oplus((ax)\gg q))\bmod b$.
Appendix B Comparison of load thresholds with achievable loads
We created datasets of random 64-bit integers and counted the number of inserted keys until the insertion fails for the first time. We tabulated the percentage of filled and empty slots in the filter to compute the achievable load of bucketed and windowed Cuckoo filters with different parameters. We show the average (main line), the minimum and maximum (shaded region) over five runs.
Figure 6: Comparison of theoretical load thresholds (horizontal blue lines), using windows and buckets and $l\in\{2,4\}$ slots per bucket or window, with achievable loads. Left: Achievable loads for different maximum random walk lengths (x-axis) for the different types of Cuckoo filters with an FPR of $2^{-14}$ and $10^{9}$ slots. Right: Achievable loads for different FPRs (color) and filter sizes (x-axis) using a fixed random walk length of 10 000.
Random walk length.
Figure 6 (left) shows the achievable loads for different maximum random walk lengths. As expected, the achievable load increases for longer random walk lengths and converges to the load threshold if enough steps are allowed for the random walk during insertion. We conclude that 10 000 steps are enough to achieve loads close to the thresholds with high probability. Note that large numbers of steps are only taken when the filter is almost at capacity; as long as it is sufficiently empty, insertion happens in a single step (or with a few steps), independently of the maximum walk length.
Achievable loads for different FPRs and filter sizes.
For a fixed random walk length, the achievable load should be independent of the FPR and the relative filter size. In practice, the achievable load appears independent of the FPR, except for small $k$ with e.g. an FPR of $2^{-4}$, for which the achieved load is lower than for higher $k$ (Figure 6 right). A possible explanation is that having only $2^{4}-1=15$ distinct fingerprints reduces the randomness for the second bucket or window. In addition, the achievable load decreases slightly for larger filters (with constant overhead factor, i.e., both $n$ and the filter size are at constant ratio), which may also be explained by the problem of mapping the small space of fingerprints to the much larger number of buckets or windows.
Appendix C Distribution of random walk lengths
Long random walks cause many cache misses and hence lead to slow insertion times. Therefore, short random walks are preferable. This can easily be achieved by increasing the space overhead of the filter, since the costly insertions happen when the filter is almost filled to its maximum capacity (see Figure 6). However, in many applications small filters are preferable, in particular, if more queries than insertions are performed.
Figure 7: Distribution of random walk lengths when inserting $10^{8}$ elements into different Cuckoo filter types for a fixed FPR of $2^{-10}$ and a filter size that ensures a load of at most $98\%$ of the theoretical load threshold.
Figure 7 shows the random walk length distribution of bucketed and windowed Cuckoo filters allowing a maximum random walk length of $10^{4}$. Most keys are directly inserted into a free slot and the filter performs a long random walk only for a very small fraction of keys; note the logarithmic scale of the y-axis. Choosing a maximum load such that the filter is only filled up to $98\%$ of its maximum capacity ensures that the actual maximum random walk length is much lower than the allowed maximum random walk length of $10^{4}$, with $(2,4)$ Cuckoo filters having shorter random walks compared to $(2,2)$ Cuckoo filters.
Appendix D Comparison of C++ and Python (with numba) implementations
To rule out that our Cuckoo filter implementation is slower or faster due to the use of just-in-time compiled Python, we compared the runtime of the original implementation of the Cuckoo filter ($(2,4)$-bucketed Cuckoo filter, computing the alternative bucket with an XOR operation, implemented in C++ [12]) with our implementation of the the same (2,4)-bucketed Cuckoo filter. Both implementations are as similar as possible, e.g., both implementations use the same trick for bit-parallel lookups, but for example, different hash functions.
Table 4: Throughput in million keys per second, single threaded; higher is better. We compare the original Cuckoo filter implementation written in C++ with an implementation in just-in-time compiled Python using numba. All filters have $2^{30}$ slots and a load of $95\%$.
FPR $2^{-5}$ (8-bit slots) FPR $2^{-13}$ (16-bit slots)
insert lookup insert lookup
C++ 7.305 61.939 7.371 60.517
numba 8.938 70.688 9.104 55.013
Table 4 shows the throughput for both implementations for Cuckoo filters configured to have a load factor of $95\%$. The throughput is averaged over three repeated runs. We measure lookup time by querying $50\%$ contained and $50\%$ not contained keys. Since both implementations have similar running times, in particular for lookups, the numba implementation does not affect our reasoning on an algorithm engineering level when comparing our novel Cuckoo filters with other state-of-the-art probabilistic filters.